i live 10 miles from work at a speed of 40 miles per hour what time should i leave home to be at work by 8:15

Answer by jim_thompson5910(35256)

d = rt

10 = 40t

10/40 = t

1/4 = t

t = 1/4

So it will take you 1/4 of an hour (or 15 minutes) to drive 10 miles at 40 mph.

So in ideal conditions, you would leave at 8:15 - 0:15 = 8:00...but, you may encounter traffic or unplanned events that will slow you down.
Also, your speed will vary and it will not be 40 mph the entire way.

So I'd say play it safe and double that 15 minute time to 2*15 = 30 minutes

So you should leave at 8:15 - 0:30 = 7:45 am

Note: it doesn't hurt to be early. So even leaving before that will work as well.